Select the correct answer.

A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?

A. -3x + 4y = 3

B. -1.5x − 3.5y = -31.5

C. 2x + y = 20

D. -2.25x + y = -9.75



1
Expert's answer
2021-07-11T18:08:59-0400

"-7x + 3y = -21.5 \/ : 3"

"\large\frac{-7}{3}x+y=-\frac{43}{6}" "\to y = \frac{7}{3}x-\frac{43}{6}"

perpendicular -> "k_1*k_2=-1"

"\large\frac{7}{3}*k_2=-1 \to k_2=-\frac{3}{7}"

B) "-1.5x-3.5y=-31.5"

"y = -\frac{3}{7}x+9"

Answer is B